# Enhancing End Stage Renal Disease Outcome Prediction: A Multi-Sourced Data-Driven Approach

### v3

@@ 6 Discussion
This study demonstrates the value of integrating multiple data sources and deep learning methodologies for improving ESRD progression predictions in CKD patients. By combining claims and EHR data with LSTM and GRU models, which excel at capturing temporal dependencies, we achieved enhanced predictive accuracy using a 24-month observation window. This window length provides an optimal balance between early detection and prediction reliability, supporting timely clinical interventions. Our research addresses a critical gap in chronic disease management by integrating clinical and claims data, offering a more comprehensive view of patient health trajectories than single-source approaches. The combination of medical indicators and healthcare utilization patterns provides deeper insights into ESRD progression risk factors, making the model more robust across diverse patient populations. The determination of this optimal observation window represents a significant advancement in balancing timely interventions with prediction accuracy, enabling healthcare providers to efficiently monitor CKD progression while minimizing data collection burden.
Our findings suggest that while the updated 2021 eGFR equation shows modest improvement in overall prediction accuracy compared to the 2009 equation (error rate reduction from 0.0232 to 0.0228) in this study population, the more useful insight may come from the pattern of misclassifications. The fact that only 7 cases were misclassified by both equations, despite similar overall error rates, indicates that these equations capture different aspects of kidney function and disease progression. This has important implications for clinical practice: while the 2021 equation may better serve minoritized populations by removing race-based adjustments, the persistence of prediction errors clustered in certain patient profiles suggests that eGFR alone may not be sufficient for accurate ESRD prediction. These findings underscore the need for a multi-faceted approach to risk assessment, combining both traditional clinical markers and novel predictive features. By leveraging integrated data sources and advanced modeling techniques, healthcare providers can develop more comprehensive risk assessment strategies, potentially leading to more targeted interventions and improved patient outcomes across diverse populations.
@@ 6.1 Limitations
Our study’s reliance on data from a single institution limits the model’s generalizability to all care delivery settings. Incorporating data from more sources could enhance robustness and applicability across different clinical settings. The framework is also currently tailored to predict ESRD outcomes specifically in CKD patients; expanding it to other chronic diseases could provide broader insights.
The use of EHR data presents challenges, such as observational bias, incomplete data, and underrepresentation of certain populations, which can impact model accuracy and fairness. Addressing these issues through comprehensive datasets or advanced imputation techniques is necessary for further improving prediction reliability.
Despite employing data oversampling methods, the disparity between our model’s high AUROC scores and relatively lower F1 and AUPRC scores indicates persistent challenges in addressing data imbalance. This gap suggests that traditional oversampling techniques may not fully compensate for the inherent imbalance in ESRD progression data. To enhance performance, exploring advanced data balancing methods is essential. Techniques such as ensemble methods, cost-sensitive learning, and hybrid approaches that combine oversampling and undersampling have shown promise in effectively managing imbalanced datasets.
Additionally, the model does not fully account for time lags between claims and clinical data, which may affect the precision of predictions. Reducing these discrepancies is essential for enhancing the model’s effectiveness in clinical practice.
@@ 6.2 Future Directions
Addressing Data Censoring: To mitigate data censoring issues, an initial strategy could involve truncating the data up to 2016 and analyzing patient outcomes in 2017 and 2018. Following this, expanding the dataset to include more recent patient information beyond 2018 would be crucial. This expansion would enhance our understanding of health trajectories and disease progression, enabling continuous refinement of the model to better align with evolving healthcare practices and patient demographics.
Integration of Unstructured Clinical Notes: Future model iterations will integrate unstructured clinical notes, such as physician narratives and discharge summaries, to capture detailed patient information often missed in structured formats. This approach aims to provide a more comprehensive view of patient health, enhancing outcome predictions and enabling more personalized interventions.
Advanced Methods for Time-Lag Mitigation: Advanced techniques to mitigate time lags between claims and clinical data, such as algorithms that adjust for expected delays or incorporate more real-time data sources for more accurate triangulation, are needed to improve the model’s ability to handle these temporal discrepancies, enhance prediction accuracy and ensure timely, data-driven interventions.
Expansion to Other Chronic Diseases: To evaluate framework’s versatility, we plan to apply it to other chronic conditions, such as heart disease. This will help assess its applicability across various care delivery settings, providing insights into its broader potential for chronic disease management and informing further optimization efforts.
@@ 7 Conclusion
As multiple sources of patient-level healthcare data continue to get integrated and analyzed in large data repositories and cloud computing platforms, this study presents an insightful illustrative approach for predicting ESRD outcome in CKD patients by integrating clinical and claims data with advanced machine learning and deep learning methods. Our results demonstrate that this combined data strategy significantly improves prediction accuracy, offering a more holistic understanding of CKD progression. Through explainable AI techniques such as SHAP analysis, we identified critical factors influencing predictions at both individual and cohort levels.
A key contribution of this work is the development of a framework that optimizes the observation window, balancing early detection with predictive accuracy to enhance clinical decision-making. The 24-month observation window was identified as the optimal period for maximizing prediction effectiveness while minimizing unnecessary interventions. Additionally, our assessment of the updated 2021 eGFR equation highlights an approach to health equity, contributing to fairer healthcare outcomes.
Overall, this study advances chronic disease management by utilizing integrated data and innovative AI techniques, paving the way for more accurate, personalized, and equitable care in CKD management. These findings lay the groundwork for future research to refine predictive models and expand their application to other chronic diseases.
@@ 8 Acknowledgments
AI Writing Assistance: All drafts of this manuscript were created by the authors without AI assistance. GPT-4o and Claude 3.5 (Sonnet) were used solely for polishing grammar and refining expression. The outputs from these tools were critically compared and reviewed by the authors to ensure accuracy and clarity. Final revisions were made by the authors to ensure all expressions align with the intended meaning and academic standards.
@@
Supplementary Table 1: Model Performance Metrics Across Observation Windows
Table 4: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data. Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).
@@
Supplementary Figure 1
Figure 7: Distribution of time to ESRD for the ESRD cohort. The plot illustrates the variability in time to ESRD across the cohort, with a mean time of 4.82 years and a standard deviation of 1.82 years.
Supplementary Table 2
Table 5: Mapping between Feature Abbreviations in The Feature Importance Plot and Full Names

### v4

@@ 6 Discussion
This study demonstrates the value of integrating multiple data sources and deep learning methodologies for improving ESRD progression predictions in CKD patients. By combining claims and EHR data with LSTM and GRU models, we achieved enhanced predictive accuracy using a 24-month observation window—an optimal balance between early detection and prediction reliability. Our approach addresses a critical gap in chronic disease management by offering a more comprehensive view of patient health trajectories than single-source approaches.
Error analysis revealed important clinical implications. Patients incorrectly predicted not to develop ESRD (false negatives) had significantly lower prevalence of CKD Stage 5 records at timestamp 6 compared to true positives (0.06 vs 0.36), suggesting sudden kidney function decline—a phenomenon clinicians recognize as ”crashing” into dialysis. Conversely, false positives showing CKD Stage 5 in consecutive timestamps without progression may reflect data censoring, as patients might have progressed after our dataset endpoint.
While the updated 2021 eGFR equation showed modest improvement in prediction accuracy compared to the 2009 equation (error rate reduction from 0.0232 to 0.0228), the pattern of misclassifications offers greater insight. With only 7 cases misclassified by both equations despite similar overall error rates, these formulas clearly capture different aspects of kidney function. This suggests that while the 2021 equation may better serve minoritized populations by removing race-based adjustments, eGFR alone remains insufficient for accurate ESRD prediction.
These findings underscore the need for a multi-faceted approach to risk assessment, combining traditional clinical markers, novel predictive features, and awareness of potential rapid disease progression patterns to develop more targeted interventions across diverse populations.
@@ 6.1 Limitations
This study’s reliance on data from a single institution may limit the model’s generalizability to other care settings. Using EHR data introduces observational bias, incomplete records, and underrepresentation of certain patient groups, which can undermine both accuracy and fairness. Although we applied oversampling to mitigate class imbalance, the contrast between high AUROC and lower F1/AUPRC indicates that imbalance remains an issue; more sophisticated approaches—such as ensemble methods, cost-sensitive learning, or hybrid sampling—may be needed. Finally, unaddressed time lags between claims and clinical data can distort temporal relationships and reduce predictive precision.
@@ 6.2 Future Directions
To address data censoring issues, we will first truncate data to 2016 and analyze outcomes from 2017-2018, then expand our dataset beyond 2018 to enhance trajectory modeling. We’ll integrate unstructured clinical notes to capture patient information missed in structured data, providing a more comprehensive health view and improving prediction accuracy.
We will implement advanced algorithms to synchronize claims and clinical data temporally, reducing time lag effects that currently impact model performance. To understand prediction errors, we’ll perform SHAP analysis on misclassified cases, identifying key features contributing to these misclassifications and guiding targeted model improvements.
Finally, we’ll validate our framework’s versatility by applying it to other chronic conditions such as heart disease, assessing its broader potential across various care delivery settings. This systematic expansion will provide actionable insights for both our ESRD prediction model and chronic disease management more broadly.
@@ 7 Conclusion
This study demonstrates the effectiveness of integrating diverse healthcare data with advanced machine learning techniques to accurately predict ESRD in CKD patients. The combined-data approach substantially enhances predictive performance and provides deeper insights into disease progression. SHAP analysis and feature importance assessment highlighted key predictors at both individual and cohort levels.
A critical contribution of this work is a framework for optimizing observation windows, balancing early detection with prediction accuracy. We identified a 24-month observation window as optimal for maximizing predictive effectiveness while minimizing unnecessary interventions. Additionally, evaluating the updated 2021 eGFR equation supports efforts towards health equity and fairer clinical outcomes.
Overall, this research advances CKD management through integrated data and innovative AI methodologies, setting the stage for personalized, equitable care and future applications in chronic disease prediction and management.
@@ 8 Acknowledgments
Contributorship Statement
Both authors meet the International Committee of Medical Journal Editors (ICMJE) authorship criteria. Rema supervised the research, advised on study conception and design, reviewed and critically revised the manuscript. Yubo contributed to study conception and design, implemented all computational methods, wrote the manuscript draft, and performed data analysis. Both authors approved the final manuscript version and agreed to be accountable for all aspects of the work, ensuring its accuracy and integrity.
AI Writing Assistance: All drafts of this manuscript were created by the authors without AI assistance. GPT-4.5 and Claude 3.7 (Sonnet) were used solely for polishing grammar and refining expression. The outputs from these tools were critically compared and reviewed by the authors to ensure accuracy and clarity. Final revisions were made by the authors to ensure all expressions align with the intended meaning and academic standards.
@@
[…]
Appendix A Detailed Data Source Information
Administrative Claims Data
The de-identified administrative claims dataset used in this study was provided by Highmark, a major health insurance provider serving Pennsylvania, Delaware, and West Virginia. The dataset includes comprehensive records of patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018. The use of this dataset was approved by Highmark’s Institutional Review Board.
Clinical EHR Data
The clinical dataset was sourced from the Electronic Health Record (EHR) data of The Nephrology Associates (TMA, P.C.), a leading community nephrology practice serving the greater Pittsburgh region. This dataset contains detailed laboratory results, patient demographics, diagnostic information, and medication records spanning from 1998 to 2018. For consistency of integration with the claims data, we truncated the clinical data to match the 10-year span of the claims data (2009–2018). The usage of the clinical data was approved by the Institutional Review Board of Carnegie Mellon University.
Appendix B Laboratory Test Value Ranges for Distribution Detection and Adjustment
Table 4: The table below presents the reference ranges established for laboratory tests during the data preprocessing phase. These ranges were used for distribution detection and adjustment to identify and correct unit inconsistencies.
Laboratory Test Minimum Value Maximum Value Unit Conversion Notes
Serum Creatinine 0 15 $1\mathrm{mg}/\mathrm{dL}=88.4\mu\mathrm{~{}mol}/\mathrm{L}$
Intact PTH 0 500 $50\mathrm{pg}/\mathrm{mL}=5.263\mathrm{pmol}/\mathrm{L}$
Hemoglobin 2 20 $6451\mathrm{~{}g}/\mathrm{dL}=10\mathrm{~{}g}/\mathrm{L};1\mathrm{~{}g}/\mathrm{dL}=100/$ 645 $\mathrm{mmol}/\mathrm{L}$
Albumin to Creatinine Ratio, Urine 0 600 $1\mathrm{mg}/\mathrm{g}=1000\mathrm{mg}/\mathrm{mg}$
Bicarbonate (Mostly Arterial) 0 30 $\mathrm{mEq}/\mathrm{L}=\mathrm{mmol}/\mathrm{L}$
Phosphorus 0 10 $4\mathrm{mg}/\mathrm{dL}=0.00129\mathrm{mmol}/\mathrm{L}$
Serum Calcium 0 20 $10\mathrm{mg}/\mathrm{dL}=0.00250\mathrm{mmol}/\mathrm{L}$
Urine Albumin 0 600 -
eGFR 0 150 -
Appendix C Comorbidity Feature ICD Code Prefixes
Table 5: Comorbidity features with ICD-9 and ICD-10 code prefixes used for feature extraction in this study
Condition ICD-9 Prefixes ICD-10 Prefixes
Diabetes 249*, 250* E08*–E13*
Hypertension 401*–405* I10*–I15*
Cardiovascular Disease 429* I25*
Anemia 285.9 D64.9, D50*–D59*
Metabolic Acidosis 276.2 E87.2
Proteinuria 791.0 R80.9
Secondary Hyperparathyroidism 252.02, 588.81 E21.1, N25.81
Phosphatemia 275.3 E83.39
Atherosclerosis 440* I70*
Congestive Heart Failure (CHF) 428* I50*
Conduction & Dysrhythmias 426*–427* I44*–I49*, Z450*, Z958*
Myocardial Infarction (MI) 410* I21*
CVD Other 390*–459* I05*–I09*, I30*–I52*
Stroke 433*–434* I63*
Fluid/Electrolytes 276* E86*–E87*
Mineral Disorders 275* E20*–E21*, E83*, N25*
Nutritional Deficiencies 260*–269* E40*–E46*, E50*–E64*, D50*
CKD Stage 4 585.4* N18.4*
CKD Stage 5 585.5* N18.5*
CKD End-Stage Renal Disease (ESRD) 585.6* N18.6*
Note: prefix * matches any sequence of trailing digits. For example, I44*–I49* captures all codes from I440 through I499 such as I4402 and I4439; Z450* captures Z4502, Z45010, Z45018, etc.
Appendix D Neural Network Implementation Details
Hardware Specifications
Resource Specification
GPU 2 $\times$ NVIDIA RTX 4090 (24GB VRAM)
RAM 512GB DDR4
Storage 2TB NVMe SSD
Table 6: Hardware specifications used for model training and evaluation.
Software Environment and Library Versions
Component Version Purpose
Python 3.10.8 Programming language
CUDA 11.6 GPU acceleration library
PyTorch 1.13.0 Deep learning framework
H2O 3.38.0.1 Machine learning model tuning
NNI 2.10 Neural architecture search
scikit-learn 1.1.3 Traditional ML algorithms
pandas 1.5.1 Data manipulation and analysis
NumPy 1.23.4 Numerical computing
Table 7: Software environment and library versions used for model implementation.
Model Architecture, Hyperparameters, and Training Details
Model Architecture Details Parameters Hyperparameters Training Time
CNN 3 conv layers (64, 128, 256 filters), 2 FC layers (512, 256) 4.3M Batch size: 32, Optimizer: Adam, Learning rate: 1e-4, Dropout: 0.3 8.5 hours
RNN 2 recurrent layers (128 units), 2 FC layers (256, 128) 2.1M Batch size: 64, Optimizer: RMSprop Learning rate: 5e-4, Dropout: 0.2 9.2 hours
LSTM 2 LSTM layers (256 units), 1 FC layer (128) 5.7M Batch size: 32, Optimizer: Adam Learning rate: 1e-4, Dropout: 0.3 15.3 hours
GRU 2 GRU layers (256 units), 1 FC layer (128) 4.9M Batch size: 32, Optimizer: Adam Learning rate: 1e-4, Dropout: 0.3 13.7 hours
TCN 4 temporal blocks, kernel size 3, 128 filters 3.8M Batch size: 32, Optimizer: Adam Learning rate: 2e-4, Dropout: 0.2 12.1 hours
Table 8: Model architecture, hyperparameters, and training details for each neural network model.
Appendix E Model Performance Metrics Across Observation Windows
Table 9: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data (24-month observation window, n=1,422). Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).
@@ Appendix E Model Performance Metrics Across Observation Windows
Appendix F Distribution of time to ESRD for the ESRD cohort
Figure 7: The plot illustrates the variability in time to ESRD across the cohort, with a mean time of 4.82 years and a standard deviation of 1.82 years.
Appendix G Feature Abbreviations
Table 10: Mapping between feature abbreviations in The feature importance plot and corresponding full names.